OpenStudy (anonymous):

how to multiply 0.024(1523.65)

OpenStudy (anonymous):

calculator

OpenStudy (anonymous):

Very, very patiently.

OpenStudy (lgbasallote):

by applying the fundamentals of multiplication

OpenStudy (anonymous):

or just multiply the value 152365*24 and then add the decimals after

OpenStudy (anonymous):

If you know basic concepts involving decimals, you have: \[10^{-3}\cdot 24\cdot 10^{-2}\cdot 152,365\]

OpenStudy (lgbasallote):

you can also use an abacus

OpenStudy (lgbasallote):

i prefer that..it's easier

OpenStudy (anonymous):

If you want to do this manually, you can do this: \[152,365\cdot 24=(100,000+50,000+2,000+300+60+5)(20+4)\\=20(100,000+50,000+2,000+300+60+5)\\+4(100,000+50,000+2,000+300+60+5)\]

OpenStudy (anonymous):

36.5676 (BE SURE TO PRESS BEST ANSWER IF I HELPED) :)

OpenStudy (anonymous):

^you just gave the answer

OpenStudy (anonymous):

Ya so.? Mind your own you just told her to use a calculator? haha

OpenStudy (anonymous):

I did not tell her to use a calculator. Your argument fails on two accounts. Try again.

OpenStudy (lgbasallote):

lol i'll medal you twice if i can @Limitless =))

OpenStudy (anonymous):

Oops my bad? you told her to be patient? whats your problem? calm down.

OpenStudy (anonymous):

No problem at all. It's just easy to point out you're violating the rules and being unproductive. See: http://openstudy.com/code-of-conduct I personally don't care about the rules, but they are rather logical.

OpenStudy (anonymous):

Okay my bad. I'll explain my answer?

OpenStudy (anonymous):

It has already been explained. If you have a method that isn't listed here, then show it.

OpenStudy (anonymous):

no its fine if your concerned, I'' be more than happy to explain myself

OpenStudy (anonymous):

I was patient! haha

OpenStudy (anonymous):

That's not really a clever or inventive method. I showed two different aspects of one way. You should try to find a method that is different from mine and the others suggested and repost an answer. Trust me, there are other methods.

OpenStudy (anonymous):

okay, okay. I'll be patient and wait for a method to come to me.........

OpenStudy (anonymous):

Are there any shortcuts to this ? I have to answer problems like those on a timed exam

OpenStudy (lgbasallote):

multiply 152365 by 24 then just put 5 decimal places

OpenStudy (lgbasallote):

fastest i can think of

OpenStudy (anonymous):

Vid, if you are having trouble with this, practice a lot of them. You will get used to it quickly. I recommend the way I suggested. It's MUCH faster than it looks.